OpenStudy (anonymous):

write an equation in point slope form of the line through point j(4,1) with the slope of -4

OpenStudy (abdullah_me):

y-1=-4(x-4)

OpenStudy (anonymous):

do you understand how Abdullah_me got that answer ?

OpenStudy (anonymous):

slope intercept form ==> y - y1 = m(x - x1) (4,1) x1 = 4 and y1 = 1 m is your slope...in this problem m = -4 now just sub y - 1 = -4(x - 4) -- you do not need to solve, just leave the parenthesis in.

OpenStudy (anonymous):

oops...typo...y - y1 = m(x - x1) is the point slope form, not slope intercept form

OpenStudy (anonymous):

Refer to the Mathematica attachment.